# Diffuson-Dominated Thermal Transport Crossover from Ordered to Liquid-like Cu3BiS3:The Negligible Role of Ion Hopping

### v1

@@ Abstract
Fundamentally understanding lattice dynamics and thermal transport behavior in liquid-like, partially occupied compounds remains a long-standing challenge in condensed matter physics. Here, we investigate the microscopic mechanisms underlying the ultralow thermal conductivity in ordered/liquid-like Cu3BiS3 by combining experimental methods with first-principles calculations. We first experimentally synthesize and characterize the ordered structure and liquid-like, partially Cu-atom occupied Cu3BiS3 structure with increasing temperature. We then combine self-consistent phonon calculations, including bubble-diagram corrections, with the Wigner transport equation, considering both phonon propagation and diffuson contributions, to evaluate the anharmonic lattice dynamics and thermal conductivity in phase-change Cu3BiS3. Our theoretical model predicts an ultralow thermal conductivity of 0.34 W·m-1·K-1 at 400 K, dominated by diffuson contributions, which accurately reproduces and explains the experimental data. Importantly, the machine-learning-based molecular dynamics (MD) simulations not only reproduced the partially Cu-atom occupied Cu3BiS3 structure with the space group but also successfully replicated the thermal conductivity obtained from experiments and Wigner transport calculations. This observation highlights the negligible impact of ionic mobility arising from partially occupied Cu sites on the thermal conductivity in diffuson-dominated thermal transport compounds. Our work not only sheds light on the minimal impact of ionic mobility on ultralow thermal conductivity in phase-change materials but also demonstrates that the Wigner transport equation accurately describes thermal transport behavior in partially occupied phases with diffuson-dominant thermal transport.
@@
American Chemical Society, LaTeXThese authors contributed equally
[…]
Thayer School of Engineering, Dartmouth College, Hanover, New Hampshire, 03755, USA
[…]
Linköping University, Department of Physics, Chemistry, and Biology (IFM), Linköping, SE-58183, Sweden
[…]
State Key Laboratory of Superhard Materials, College of Physics, Jilin University, Changchun, 130012, China
Keywords: thermoelectric material, Cu-based chalcogenide, lattice anharmonicity, wigner transport equation, self-consistent phonon theory
[…]
Some journals require a graphical entry for the Table of Contents. This should be laid out “print ready” so that the sizing of the text is correct.
Inside the tocentry environment, the font used is Helvetica 8 pt, as required by Journal of the American Chemical Society.
The surrounding frame is 9 cm by 3.5 cm, which is the maximum permitted for Journal of the American Chemical Society graphical table of content entries. The box will not resize if the content is too big: instead it will overflow the edge of the box.
This box and the associated title will always be printed on a separate page at the end of the document.
1 Introduction
@@
Ultra-low lattice thermal conductivity () in solid-state materials is crucial for advancing energy conversion technologies 1, 2, improving fuel cell performance 3, and optimizing various thermoelectric applications in the fields of thermal barrier coatings 4, waste heat recovery 5, 6. However, understanding and measuring lattice thermal conductivity is particularly challenging for materials undergoing phase transitions 7. These materials are widely available and have been extensively used in various fields, including thermoelectrics 8, solid-state memory 9, and switches 10. Therefore, a fundamental understanding of the origins of ultra-low thermal conductivity through a comprehensive investigation of the relationships between transport phenomena, lattice dynamics, and structural configurations is of great significance. This understanding unveils critical insights and establishes foundational principles for the innovation and discovery of advanced material systems 11.
@@
Within the family of phase transition materials, Cu-based chalcogenides stand out due to their complex phase transition behaviors and outstanding thermoelectric properties 12, 13. Cu-based chalcogenides often undergo complex lattice distortions and significant changes in thermal transport properties during phase transitions, making them a key platform for exploring new thermal conduction mechanisms. For instance, Liu et al. 14 reported a significant reduction in the thermal conductivity of Cu2Se, attributed to its second-order phase transition, which led to a notable enhancement of the thermoelectric figure of merit () from 0.6 to 2.3. The recent synthesis of copper-rich wittichenite Cu3BiS3 15, 16, which exhibits intrinsically low lattice thermal conductivity, is particularly noteworthy. This material provides a compelling platform to investigate structural features—such as atomic coordination and cationic networks—that influence thermal transport. Meanwhile, the present investigation on Cu3BiS3 clearly demonstrates that the 3-fold coordination of copper, combined with its strong anisotropic vibrations, serves as the driving force behind its ultra-low thermal conductivity 17. Despite extensive research on Cu3BiS3, its phase transition behavior remains unclear, and the mechanisms underlying its ultra-low thermal conductivity across different phases have yet to be fully elucidated.
@@
Historically, thermal transport in crystalline materials has been well described by the phonon-gas model, which treats phonons as quasiparticles propagating over long distances within the framework of the conventional phonon Boltzmann transport equation (BTE) 18, 19. Within the conventional BTE framework, point defects 20, anharmonicity 21, 22, and bond strength variations 23, 24 act as phonon scattering sources, reducing mean free paths and thereby lowering lattice thermal conductivity. However, certain complex crystalline materials exhibit ultra-low thermal conductivities that cannot be fully explained by the phonon-gas model 25. In such cases, phonon mean free paths are extremely short—comparable to or even smaller than interatomic distances—resulting in the loss of their quasi-particle-like nature 26. To address the limitations of the conventional BTE, Simoncelli et al. 27 developed a unified theory of thermal transport by reformulating the Boltzmann Transport Equation within the Wigner formalism. Specifically, the Wigner transport equation 28 describes thermal transport via the phonon velocity operator, wherein the diagonal and off-diagonal elements capture both the particle-like propagation of phonons and the wave-like tunneling between eigenstates. The wave-like tunneling arises from the overlap of phonon modes with similar frequencies—substantially broadened by anharmonicity—thereby facilitating thermal energy transfer between modes. The unified theory of thermal transport has been successfully applied to address the underestimation of thermal conductivity within the BTE framework for strongly anharmonic systems, such as Cu12SbS29, La2Zr2O727, Cs2AgBiBr6 30.
@@
Another challenge in evaluating the thermal conductivity of phase-change materials is the potential occurrence of a partial-crystalline partial-liquid state (PCPLS), driven by the liquid-like migration of Cu ions, as observed in superionic systems such as Ag9GaSe631, AgCrSe232, and Li2S 33. It is also reported that the most representative -Cu2S, with space group and Pearson symbol cF208, exhibits a remarkably high degree of disorder, with 8 Cu ions distributed over 204 possible sites within the unit cell 34, 35. The strikingly structural disorder suggests an exceptionally low activation barrier for Cu-ion migration, giving rise to liquid-like cationic dynamics. This remarkable ion mobility facilitates strong phonon scattering mechanisms, thereby reducing the specific heat via the suppression of transverse phonon modes 36. In contrast, Bernges et al. 37 conducted the first direct comparison of thermal and ionic transport in nine Ag+ argyrodite compositions. Their results revealed that, despite the ionic conductivity varying by several orders of magnitude, the thermal conductivity remains largely constant, indicating that high ionic mobility is not essential for diffusion-mediated thermal transport. Similarly, Ghata et al. 13 observed that with increasing temperature, the ionic conductivity of Cu7PSe6 rises significantly, while the thermal conductivity remains nearly constant, suggesting a negligible influence of ionic transport on thermal conductivity. Therefore, a fundamental understanding of the thermal transport mechanism in phase-transition and partially occupied materials remains an open question and an urgent need.
@@
In this work, we systematically investigate the anharmonic lattice dynamics and thermal transport behavior in ordered and elemental-partially occupied, liquid-like Cu3BiS3 by combining experimental and theoretical approaches. Experimentally, we synthesize and characterize the ordered and liquid-like phases of Cu3BiS3, which adopt the space groups and , respectively. We then combine first-principles calculations with machine-learning-based molecular dynamics simulations to accurately reproduce the partially Cu-atom occupied phase of Cu3BiS3 () at high temperatures. Subsequently, we applied the state-of-the-art Wigner transport equation, accounting for both propagative and diffuson contributions, to calculate the thermal conductivity of the ordered and phases of Cu3BiS3, achieving excellent agreement with experimental results. To investigate the impact of ion hopping on thermal transport at high temperatures in Cu3BiS3, we calculate the lattice thermal conductivity () using the machine-learning-based Green–Kubo equilibrium molecular dynamics (GK-EMD) method 38, which explicitly accounts for all contributions to the heat-flux operator 39, 40. We find that the predicted thermal conductivity accurately reproduces both the experimental data and the results obtained from the Wigner transport equation, indicating a negligible effect of ion mobility on thermal transport. Finally, we explore the microscopic origins of the ultralow thermal conductivity observed in both phases of Cu3BiS3. Our work not only highlights the negligible role of ionic mobility in thermal transport but also provides deeper insights into the microscopic mechanisms underlying the ultralow thermal conductivity observed in ordered and liquid-like Cu3BiS3 and related compounds.
2 Results and discussion
2.1 Experimental crystal structures
@@
Rietveld refinement of the powder X-ray diffraction (PXRD) data, carried out using the GSAS-II software package 41, 42, shows excellent agreement with the calculated patterns (R = 7.86%, R = 10.22%, and = 1.12), as illustrated in Figure 1(a). As evidenced by the well-defined PXRD patterns at 300 K, the compound is accurately indexed to the non-centrosymmetric space group , consistent with previously reported results 17, 43. Specifically, the structure comprises three Cu atoms, three S atoms, and one Bi atom, all positioned at Wyckoff 4a sites with a site multiplicity of 4. Each Bi atom is bonded to three S atoms, forming a [BiS3]3- trigonal pyramid that contains a lone pair of electrons. In contrast, each Cu atom is coordinated by three S atoms, forming a distorted trigonal planar arrangement. The unit cell parameters determined from Rietveld refinement of the phase-pure sample, namely, a = 7.6750(2) Å, b = 10.4083(3) Å, and c = 6.7060(2) Å, are in good agreement with density functional theory (DFT) calculations performed using the PBEsol functional 44 (a = 7.4618(8) Å, b = 10.4705(9) Å, c = 6.3209(5) Å) as well as with previously reported results 16, 17.
Upon further increasing the temperature, we observe a notable change in the XRD pattern (see Figure S1 in the Supporting Information). Based on the structural refinments of the high-T PXRD pattern, we identify an experimental phase transition from the orthorhombic space group to the orthorhombic space group is observed, as illustrated in Figure 1(b). This finding diverges from the previously proposed space group for CuBiS2 reported by Wu et al.in Ref.16. Although our results differ from those reported by Wu et al., it is noteworthy that the isostructural space Cu3SbS3 has been experimentally observed to undergo a phase transition from the (-) phase to the (-) phase 45, 46. This observation aligns with our findings, supporting the validity of the observed phase transition in Cu3BiS3. Interestingly, the distribution of disordered Cu+ ions within the three-dimensional framework of [BiS3]3- units has been elucidated in the phase, matching that observed in -Cu3SbS3 47. Specifically, single-crystal data reveal that Cu+ ions exhibit disorder, being partially distributed over five distinct sites, resulting in either trigonal-planar or tetrahedral coordination. To investigate this phenomenon, we conducted 20 ps ab initio molecular dynamics (AIMD) simulations at 500 K (see Figure 1(c) and Figure S2 in the Supporting Information). Notably, the simulations successfully captured the diverse sites occupied by Cu+ ions, confirming their mobility and aligning closely with our experimental observations.
@@
To elucidate the chemical environment of Cu+ ions, we performed electronic structure calculations on the phase, examining both bonding and antibonding states to gain deeper insights into its bonding characteristics. In the phase, various Cu–S bonds exhibit distinct bonding characteristics (see Figure 1(d)); nevertheless, these bonds generally share common features. All Cu–Se bonds exhibit occupied antibonding states below the Fermi level, contributing approximately 45% to the overall antibonding character. Previous studies have attributed this behavior to the proximity of the Cu d-orbitals to the energy levels of the S p-orbitals, leading to strong p–d hybridization 48. Such robust hybridization pushes the occupied antibonding states closer to the Fermi level, weakening the chemical bonds and promoting metavalent-like behavior 24, 49. The combination of low integrated electron population and substantial antibonding contributions highlights the intrinsically weak bonding environment of Cu+ within the structure, thus facilitating rapid Cu+ ion transport 50. Furthermore, the significant presence of occupied antibonding states below the Fermi level notably enhances lattice anharmonicity, substantially influencing thermal transport, as discussed in subsequent sections. Also, such a weak bonding leads to larger mean square displacement (MSD) of Cu atoms compared to the other atoms (see Figure S3 in the Supporting Information). This is consistent with the experimental observation of the atomic displacement parameters of Cu3BiS3 atoms 17.
2.2 Diffuson-dominanted thermal transport

### v2

@@ Abstract
Fundamentally understanding lattice dynamics and thermal transport behavior in liquid-like, partially occupied compounds remains a long-standing challenge in condensed matter physics. Here, we investigate the microscopic mechanisms underlying the ultralow thermal conductivity in ordered/liquid-like Cu3BiS3 by combining experimental methods with first-principles calculations. We first experimentally synthesize and characterize the ordered structure and liquid-like, partially Cu-atom occupied Cu3BiS3 structure with increasing temperature. We then combine self-consistent phonon calculations, including bubble-diagram corrections, with the Wigner transport equation, considering both phonon propagation and diffuson contributions, to evaluate the anharmonic lattice dynamics and thermal conductivity in phase-change Cu3BiS3. Our theoretical model predicts an ultralow thermal conductivity of 0.34 W·m-1·K-1 at 400 K, dominated by diffuson contributions, which accurately reproduces and explains the experimental data. Importantly, the machine-learning-based molecular dynamics (MD) simulations not only reproduced the partially Cu-atom occupied Cu3BiS3 structure with the space group but also successfully replicated the thermal conductivity obtained from experiments and Wigner transport calculations. This observation highlights the negligible impact of ionic mobility arising from partially occupied Cu sites on the thermal conductivity in diffuson-dominated thermal transport compounds. Our work sheds light on the minimal impact of ionic mobility on ultralow thermal conductivity in phase-change materials. It demonstrates that the Wigner transport equation accurately describes thermal transport behavior in partially occupied phases with diffuson-dominant thermal transport.
@@
Ion mobility, ultra-low thermal conductivity, diffuson-dominated thermal transport, phase transition, machine-learning molecular dynamics simulation, Green-Kubo equation, Wigner transport equation1 Introduction
@@
Ultra-low lattice thermal conductivity () in solid-state materials is crucial for advancing energy conversion technologies [1, 2], improving fuel cell performance [3], and optimizing various thermoelectric applications in the fields of thermal barrier coatings [4], waste heat recovery [5, 6]. However, understanding and measuring lattice thermal conductivity is particularly challenging for materials undergoing phase transitions [7]. These materials are widely available and have been extensively used in various fields, including thermoelectrics [8], solid-state memory [9], and switches [10]. Therefore, a fundamental understanding of the origins of ultra-low thermal conductivity through a comprehensive investigation of the relationships between transport phenomena, lattice dynamics, and structural configurations is of great significance. This understanding unveils critical insights and establishes foundational principles for the innovation and discovery of advanced material systems [11].
@@
Within the family of phase transition materials, Cu-based chalcogenides stand out due to their complex phase transition behaviors and outstanding thermoelectric properties [12, 13]. Cu-based chalcogenides often undergo complex lattice distortions and significant changes in thermal transport properties during phase transitions, making them a key platform for exploring new thermal conduction mechanisms. For instance, Liu et al. [14] reported a significant reduction in the thermal conductivity of Cu2Se, attributed to its second-order phase transition, which led to a notable enhancement of the thermoelectric figure of merit () from 0.6 to 2.3. The recent synthesis of copper-rich wittichenite Cu3BiS3 [15, 16, 17], which exhibits intrinsically low lattice thermal conductivity, is particularly noteworthy. This material provides a compelling platform to investigate structural features—such as atomic coordination and cationic networks—that influence thermal transport. Meanwhile, Maji et al. demonstrate that the 3-fold coordination of copper, combined with its strong anisotropic vibrations, serves as the driving force behind the ultra-low thermal conductivity in Cu3BiS3 [17]. Despite extensive researches on Cu3BiS3, its phase transition behavior remains unclear, and the mechanisms underlying its ultra-low thermal conductivity across different phases have yet to be fully elucidated.
@@
Historically, thermal transport in crystalline materials has been well described by the phonon-gas model, which treats phonons as quasiparticles propagating over long distances within the framework of the conventional Boltzmann transport equation (BTE) [18, 19]. Within the conventional BTE framework, point defects [20], anharmonicity [21, 22], and bond strength variations [23, 24] act as phonon scattering sources, reducing mean free paths and thereby lowering lattice thermal conductivity. However, certain complex crystalline materials exhibit ultra-low thermal conductivities that cannot be fully explained by the phonon-gas model [25]. The conventional BTE assumes well-separated phonon branches, and it therefore fails when this separation collapses, or when off-diagonal velocity terms become non-negligible—situations common in strongly anharmonic or disordered materials [26, 27]. To address the limitations of the conventional BTE, Simoncelli et al. [28] developed a unified theory of thermal transport by reformulating the Boltzmann Transport Equation within the Wigner formalism. Specifically, the Wigner transport equation [29] describes thermal transport via the phonon velocity operator, wherein the diagonal and off-diagonal elements (diffuson described by Allen and Feldman [26, 30]) capture both the particle-like propagation of phonons and the wave-like tunneling between eigenstates. The wave-like tunneling arises from the overlap of phonon modes with similar frequencies—substantially broadened by anharmonicity—thereby facilitating thermal energy transfer between modes. The Wigner transport equation has been successfully applied to address the underestimation of thermal conductivity within the BTE framework for strongly anharmonic systems, such as Cu12Sb4S13[31], La2Zr2O7[28], Cs2AgBiBr6 [32].
@@
Another challenge in evaluating the thermal conductivity of phase-change materials is the potential occurrence of a partial-crystalline partial-liquid state (PCPLS), driven by the liquid-like migration of Cu ions, as observed in superionic systems such as Ag9GaSe6[33], AgCrSe2[34], and Li2S [35]. It is also reported that the most representative -Cu2S, with space group and Pearson symbol cF208, exhibits a remarkably high degree of disorder, with 8 Cu ions distributed over 204 possible sites within the unit cell [36, 37]. The strikingly structural disorder suggests an exceptionally low activation barrier for Cu-ion migration, giving rise to liquid-like cationic dynamics. This remarkable ion mobility facilitates strong phonon scattering mechanisms, thereby reducing the specific heat via the suppression of transverse phonon modes [38]. In contrast, Bernges et al. [39] conducted the first direct comparison of thermal and ionic transport in nine Ag+ argyrodite compositions. Their results revealed that, despite the ionic conductivity varying by several orders of magnitude, the thermal conductivity remains largely constant, indicating that high ionic mobility is not essential for diffuson-mediated thermal transport. Similarly, Ghata et al. [13] observed that with increasing temperature, the ionic conductivity of Cu7PSe6 rises significantly, while the thermal conductivity remains nearly constant, suggesting a negligible influence of ionic transport on thermal conductivity. Therefore, a fundamental understanding of the thermal transport mechanism in phase-transition and partially occupied materials remains an open question and an urgent need.
@@
In this work, we systematically investigate the anharmonic lattice dynamics and thermal transport behavior in ordered and elemental-partially occupied, liquid-like Cu3BiS3 by combining experimental and theoretical approaches. Experimentally, we synthesize and characterize the ordered and liquid-like phases of Cu3BiS3, which adopt the space groups and , respectively. We then combine first-principles calculations with machine-learning-based molecular dynamics simulations to accurately reproduce the partially Cu-atom occupied phase of Cu3BiS3 () at high temperatures. Subsequently, we applied the state-of-the-art Wigner transport equation, accounting for both propagative and diffuson contributions, to calculate the thermal conductivity of the and phases of Cu3BiS3, achieving excellent agreement with experimental results. To investigate the impact of ion hopping on thermal transport at high temperatures in Cu3BiS3, we calculate the lattice thermal conductivity () using the machine-learning-based Green–Kubo equilibrium molecular dynamics (GK-EMD) method [40], which explicitly accounts for all contributions to the heat-flux operator [41, 42]. We find that the predicted thermal conductivity accurately reproduces both the experimental data and the results obtained from the Wigner transport equation, indicating a negligible effect of ion mobility on thermal transport. Finally, we explore the microscopic origins of the ultralow thermal conductivity observed in both phases of Cu3BiS3. Our work not only highlights the negligible role of ionic mobility in thermal transport but also provides deeper insights into the microscopic mechanisms underlying the ultralow thermal conductivity observed in ordered and liquid-like Cu3BiS3 and related compounds.
2 Experimental crystal structures
@@
Rietveld refinement carried out using the Fullprof software package [43, 44], leads to calculated patterns in excellent agreement with the powder X-ray diffraction (PXRD) data, as illustrated in Figure 1. As evidenced by the well-defined refinement at 300 K (Rp = 7.96, Rwp = 10.3, Rexp = 9.13 and = 1.28), the PXRD pattern is accurately indexed to the non-centrosymmetric space group , consistent with previously reported results [17, 45, 46]. Specifically, the crystal structure is characterized by Cu, Bi and S atoms located on three, one and three differents crystallographic sites, respectively, of general position, i.e. Wyckoff 4a. Each Bi atom is bonded to three S atoms at distances ranging between 2.565 Å and 2.614 Å (average Bi-S distance of 2.594 Å), forming a [BiS3]3- trigonal pyramid. Note that the lone pair (LP) of electrons of Bi3+ cation very probably constitutes the missing apex of a tetrahedron [BiS3LP]3-. In contrast, each Cu atom is coordinated by three S atoms at distances ranging between 2.178 Å and 2.381 Å (average Cu-S distance of 2.285 Å), forming a distorted trigonal planar arrangement. The unit cell parameters determined from Rietveld refinement of the phase-pure sample, namely, a = 7.6769(2) Å, b = 10.4107(3) Å, and c = 6.7076(2) Å, are in agreement with density functional theory (DFT) calculations performed using the PBEsol functional [47] (a = 7.4618(8) Å, b = 10.4705(9) Å, c = 6.3209(5) Å) as well as with previously reported results [16, 17].
Upon further increasing the temperature, we observe a notable change in the XRD pattern (see Figure S1 in the Supporting Information) confirming previous results reported by Makovicky highlighting that Cu3BiS3 exhibits a structural transition from a RT orthorhombic () structure to an intermediate modulated structure at 391.5 K and, above 463 K, a transition to a high temperature (HT) polymorph () [46]. Based on the structural refinments of the high-T PXRD patterns, we confirm an experimental phase transition from the orthorhombic space group to the orthorhombic space group , as illustrated in Figure 1, even if the space group is not excluded [46]. This space group change from at 300 K to at 500 K is related to the extinction of the (hk0) reflections for which h = 2n+1, as exemplified with the (120) and (310) reflections observed at 20.6° and 36.1° in 2 on the PXRD pattern recorded at 300 K and absent at 500 K. In addition, a broad peak observed near 2 13.2° does not match the known reflections of the Cu3BiS3 phase and is significantly wider than the main phase peaks. This suggests that it may arise from the sample environment or from a poorly crystallized impurity phase that could not be reliably identified. On the other hand, we have observed that CuBiS2 exists in the crystal system based on the structural data reported by Ref. [48]. To rule out the possibility of CuBiS2 as a minority or mis-indexed phase, we simulated powder-diffraction patterns for the three most plausible structures at the Cu, K wavelength: (i) Cu3BiS3 in and , the two space groups obtained from our Rietveld refinements, and (ii) CuBiS2 in , using the atomic coordinates reported by Portheine and Nowacki [48] (see Figure S2 in the Supporting Information). The calculated pattern for CuBiS2 displays an intense reflection at 2 12°, a feature that is completely absent in our experimental data, where the first detectable peak appears above 15°. Additional reflections predicted for CuBiS2 (e.g., 2 18.3°, 26.9°, 31.7°) are also missing or mismatched in both position and intensity. These systematic discrepancies unambiguously exclude CuBiS2 as a constituent phase in the investigated sample. To quantify the competition between ordered and liquid-like states from first principles, we have constructed a Cu–Bi–S phase diagram based on the convex-hull of DFT formation enthalpies (see Figure S3(a) in the Supporting Information). The diagram encompasses all computed compositions and explicitly resolves the two Cu3BiS3 polymorphs: the orthorhombic phase and the ordered variant. Notably, the structure lies 3.16 meV/atom below , i.e. = –3.16 meV/atom, confirming its thermodynamic preference.
@@
Interestingly, the distribution of Cu+ ions within the framework of [BiS3]3- units has been elucidated in the phase (see Figure S3(b) in the Supporting Information), matching that observed in -Cu3SbS3 [49]. Specifically, single-crystal data recorded at 493 K reveal that Cu+ ions exhibit disorder (see Figure S3(c) in the Supporting Information), being distributed over five distinct sites each being partially occupied, resulting in either trigonal-planar or tetrahedral coordination. To investigate this phenomenon, we conducted 20 ps ab initio molecular dynamics (AIMD) simulations at 400 K and 500 K (see Figure 2(a) and Figure S3(d-e) in the Supporting Information). Notably, the simulations successfully captured the diverse sites occupied by Cu+ ions, confirming their mobility and aligning closely with our experimental observations. To elucidate the chemical environment of Cu+ ions, we performed electronic structure calculations on the phase, examining both bonding and antibonding states to gain deeper insights into its bonding characteristics. In the phase, various Cu–S bonds exhibit distinct bonding characteristics (see Figure 2(b)); nevertheless, these bonds generally share common features. All Cu–S bonds exhibit occupied antibonding states below the Fermi level, contributing approximately 45% to the overall antibonding character. Previous studies have attributed this behavior to the proximity of the Cu d-orbitals to the energy levels of the S p-orbitals, leading to strong p–d hybridization [50]. Such robust hybridization pushes the occupied antibonding states closer to the Fermi level, weakening the chemical bonds and promoting metavalent-like behavior [24, 51]. The combination of low integrated electron population and substantial antibonding contributions highlights the intrinsically weak bonding environment of Cu+ within the structure, thus facilitating rapid Cu+ ion transport [52]. Furthermore, the significant presence of occupied antibonding states below the Fermi level notably enhances lattice anharmonicity, substantially influencing thermal transport, as discussed in subsequent sections. Also, such a weak bonding leads to larger mean square displacement (MSD) of Cu atoms compared to the other atoms (see Figure S4 in the Supporting Information). This is consistent with the experimental observation of the atomic displacement parameters of Cu atoms reported in Cu3BiS3 [17].
3 Diffuson-dominated thermal transport